A negative charge q = -2.80 Times 10^-6 C is locat
Answer #1

magnetic field due to a moving charge is B = (mu_o/4pi)*q*(v X r)/r^2

mu_0 = 4*pi*10^-7


mu_o/(4*pi) = 10^-7

v X r = [(7.5*10^4) i + (-4.9*10^4) j] X (0.14 i - 0.28 j)

v X r = -2.1*10^4 k + 0.686*10^4 k

v X r = -1.414*10^4 k

r^2 = x^2+y^2 = 0.14^2+0.28^2 = 0.098 m^2

B = (mu_o/4pi)*q*(v X r)/r^2

B = (10^-7*1.414*10^4*1.6*10^-19)/0.098 k


B=2.308*10^-21 T = 2.308*10^-15 uT k


therefore Bx = 0 T

By = 0 T

Bz = 2.308*10^-15 uT